Question 4 - Supply and Demand

Assume that the demand and the supply in a market are represented by the following equations:

Qd = 200- 3P

Qs = 2P - 100

(i) Compute and illustrate (i.e. show with a graph – make sure it is labeled) the market equilibrium in this case.

(ii) Find the CS and PS in the market and interpret what the values mean.
Answer #1

Answer:

Given,

Qd = 200- 3P

Qs = 2P - 100

(i) Thus the condition of market equilibrium is Qd=Qs=>200- 3P=2P - 100=>5P=300=>P=60

Hence the equilibrium price, P=60 and the equilibrium quantity, Q=200- 3P=200-3*60=200-180=20

Market Equilibrium Price -100 -80 -60 -40 -20 0 20 40 60 80 100 120 140 160 180 200 220 Quantity -Supply Curve +Demand Curve

(ii) CS and PS:

CS=0.5*(66.67-60)*20=66.7

and PS=0.5*(60-50)*20=100

Meaning: CS=66.7 indicates that at price=60 the consumer buys 20 quantity for which he/she is willing and can pay more up to the amount of 66.7. Similarly, PS=100 indicates that the producer also in surplus of 100 selling at price of 60 i.e. the producer is gaining more than he/she is actually want.